OpenStudy (anonymous):

x + y = 4 x - 2y = -5 If the first equation is multiplied by 2 and then the equations are added, the result is _____.

OpenStudy (anonymous):

x + y = 4 multiplied by 2 gives 2x + 2y = 8 x - 2y = -5 ------------------ 3x + 0=3 x=1

OpenStudy (jackellyn):

2(x+y)=2(4) 2x + 2y = 8 2x + 2y =8 + x - 2y =-5 ------------- 3x + 0 = 3 So 3x = 3

OpenStudy (anonymous):

Thanks again. Just checking my work